A wind turbine has 12,000 kgkg blades that are 40 mm long. The blades spin at 24 rpmrpm .

You may want to review (Page) .

Part A

If we model a blade as a point mass at the midpoint of the blade, what is the inward force necessary to provide each blade's centripetal acceleration?
Answer #1

Centripetal force required  F = mv2/

Where m is the mass

v is the velocity and

r is the distance of blade from centre

V=rW and w=21V

\omega= 2.512 radian/sec

v = 40*2.512= 100.48m/s

F=3.03×106 N

Like